# neurips_data_2024.bib

### v1

@@ 3 From link prediction to link forecasting
In Figure 4(a) we use normalized mutual information (NMI) [cover2006elements] to measure the information loss caused by splitting temporal edges $E$ into batches. NMI quantifies how much information observing one random variable conveys about another random variable. It takes values between 00, meaning “no information”, and $1$, meaning “full information”. By treating the index $i$ of each batch $B_{i}$ assigned to each edge $(u,v)\in B_{i}$ as one random variable and the associated edge’s timestamp $t$ as the other, we can measure the temporal information that is retained after dividing edges into batches. In this case, a value of $1$ means that we can reconstruct the timestamps of edges correctly from their batch number, and a value of 00 means that batch numbers do not carry any information about timestamps. Consequently, small NMI values indicate a large loss of temporal information due to batching.
@@
In Figure 4(a) we see that in continuous-time temporal graphs where timestamps have a high resolution, larger batches result in larger information loss because assigning edges that occur at different times to the same batch discards their temporal ordering; the larger the batch size, the more information is lost. A batch size of $b=1$ preserves all temporal information – i.e. maximum NMI – because we obtain a bijective mapping between timestamps and batch numbers. An exception is, e.g., the Enron dataset where emails sent to multiple recipients are recorded as simultaneously occurring edges. This leads to a maximum NMI roughly at a batch size that equals the average number of edges per snapshot.
(a) For continuous-time temporal graphs, the temporal ordering of edges within batches is discarded. With increasing batch size, more edges with different timestamps are assigned to the same time window, thus, losing more information.
(b) Small batches for discrete-time temporal graphs implicitly define an edge ordering within snapshots that is not present in the data, consequently losing the information that edges of the same snapshot occur at the same time. The NMI has its maximum near the average snapshot size (refer to Table 1) after which the values decrease again similar to continuous-time datasets.
Figure 4: Temporal information loss in terms of Normalized Mutual Information (NMI, y-axis) for different batch sizes (x-axis), where smaller NMI scores indicate more information loss.
@@
Figure 4(b) shows the batch-size dependent NMI for discrete-time temporal graphs. Similar to the Enron dataset, the “optimal” batch size that retains most temporal information depends on the average number of links per snapshot and, thus, on the characteristics of the data. Too small batch sizes impose an ordering on the edges within the snapshots that is not present in the data while too large batches stretch across snapshots and discard the temporal ordering of edges from different batches.
These results show that changing the batch size influences both the prediction time window as well as the temporal information available to TGNNs. Effectively, the batch size is a hidden hyperparameter that directly impacts the characteristics (and difficulty) of the prediction task. In real-world applications, however, the prediction time window is inherently connected to the problem at hand, necessitating a task formulation that is chosen carefully for each benchmark instead of for each model. To address this issue, we propose a new task formulation that utilizes a fixed prediction time window thereby solving the first problem. Using a fixed time window rather than a fixed number of links also mitigates the second problem by providing an equal amount of temporal information in each batch, thus facilitating a fair model comparison.
Link forecasting: task definition.
@@
The study of temporal information is at the center of time series forecasting and, therefore, we relate our task definition to a fixed temporal quantity to solve the identified problems [DeepLearningTime2023benidis]. We can interpret the temporal edges $E$ as $n^{2}$ Boolean time series, each of which takes the value 1 at those times when an edge occurs. Standard multivariate models output a value for each timestamp over a forecasting horizon $h$. In large-scale temporal graphs, it is computationally infeasible to forecast the existence of all $n^{2}$ possible links, thus, only a sample of negative edges is considered instead. In continuous-time dynamic graphs, observations are available at high resolution, e.g. seconds, however, for many practical applications, predicting at lower granularity suffices. For example, it is typically enough to predict whether a customer purchases a certain product within the next day or week. Therefore, we consider forecasting for all timestamps $[t+1,t+h]$ at once instead of for each of them individually, and define the link forecasting task as follows:
Definition 3.1 (Dynamic link forecasting).
@@
Let $G=(V,E)$ be a temporal graph with node features $H_{V}$ and edge features $H_{E}$. Let $h$ be the time horizon and $W_{i}^{+}:=\{(u,v)\mid\exists(u,v,t)\in E\text{ with }i\cdot h\leq t<(i+1)\cdot h\}$ the set of edges in the $i$-th time window. We use $W_{i}^{-}$ to denote a sample of negative edges that do not occur in time window $[i\cdot h,(i+1)\cdot h)$. We further use $\hat{E}_{i}=\left\{(u,v,t)\mid\exists(u,v,t)\in E:t<i\cdot h\right\}$ to denote the set of past edges for time window $i$. The goal of dynamic link forecasting is to find a model $f_{\theta}(u,v|\hat{E}_{i},H_{V},H_{\hat{E}_{i}})$ with parameters $\theta$ that, for each time window $i$ forecasts whether $(u,v)\in W_{i}^{+}$ or $(u,v)\in W_{i}^{-}$.
Crucially, this definition makes the evaluation independent of the batch size $b$ and instead introduces a time horizon $h$ that defines the forecasting time window. Note that we define the past edges $\hat{E}$ as a temporal graph to keep temporal information, e.g., to enable sampling of recent neighbors [TemporalGraphNetworks2020rossi]. The forecasting time window, on the other hand, is a snapshot that discards temporal information inside each window since we are forecasting if a link exists in any of the timestamps $t\in[t+1,t+h]$. For parallel processing, batches can be defined either as $E_{t+1:t+h}\cup E^{-}_{t+1:t+h}$ and can have different sizes, or as a subset $B\subset E_{t+1:t+h}\cup E^{-}_{t+1:t+h}$ as long other time-dependent parts of the optimization are based on $E_{t+1:t+h}$, e.g., negative sampling or updating a memory.
@@
We provide implementations for our evaluation procedure in commonly used PyTorch [PyTorchImperativeStyle2019paszke] libraries to simplify the adoption of our approach. Specifically, we implement a new DataLoader called SnapshotLoader [^1] https://github.com/M-Lampert/pytorch_geometric/blob/snapshot-loader/torch_geometric/loader/snapshot_loader.py that replaces the widely used TemporalDataLoader in PyTorch Geometric [FastGraphRepresentation2019fey]. We extend DyGLib [BetterDynamicGraph2023yu] with a command line argument horizon that can be used in the evaluation pipeline. [^2] https://github.com/M-Lampert/DyGLib The latter was used for the experiments in this work and can be used to reproduce our results.
4 Link prediction vs. forecasting in state-of-the-art TGNNs
@@
We now experimentally evaluate nine state-of-the-art models [PredictingDynamicEmbedding2019kumar, DyRepLearningRepresentations2019trivedi, InductiveRepresentationLearning2020xu, TemporalGraphNetworks2020rossi, InductiveRepresentationLearning2021wang, BetterEvaluationDynamic2022poursafaei, TCLTransformerbasedDynamic2021wanga, WeReallyNeed2023cong, BetterDynamicGraph2023yu], evaluating their performance both based for the (usual) dynamic link prediction and the proposed dynamic link forecasting task. We use implementation and model configurations provided by DyGLib [BetterDynamicGraph2023yu] (cf. Appendix C) and repeat each experiment five times to obtain averages. We use historical negative sampling [BetterEvaluationDynamic2022poursafaei] and train each model using batch-based training and validation with batch size $b=200$ which was found “to be a good trade-off between speed and update granularity” [TemporalGraphNetworks2020rossi] and adopted in similar works [BetterDynamicGraph2023yu, BetterEvaluationDynamic2022poursafaei]. Afterwards, we evaluate each model trained with batch-based training using our proposed time-window-based evaluation method as well as the common batch-based evaluation approach with $b=200$. We choose the forecasting horizon $h$ such that we obtain average batch sizes of approximately $200$ for all continuous-time datasets to make the results of our new evaluation method comparable to the results of the batch-oriented approach (Table 2). For discrete-time temporal graphs, we set $h=1$ to obtain one time window per snapshot, predicting for time intervals ranging from five minutes to a year, depending on the datasets (Table 1).
@@
Table 2: Average links per window $|E_{t_{i}:t_{j}}|$ and standard deviation for horizon $h$ used in the evaluation (left). We chose $h$ for each dataset to get $\overline{|E_{t_{i}:t_{j}}|}\approx 200$. The average time window duration is in hours and seconds per batch (center). NMI uses the window and batch IDs of the test set as random variables quantifying how much the chosen chunks differ between the approaches (right).
Dataset $h$ $\overline{|E_{t_{i}:t_{j}|}}$ $b$ Avg Duration (h) Avg Duration (s) NMI
Enron 172 800s ($48$h) 214.1 ± 274.1 200 50.1 ± 165.38 180395.2 ± 595358.23 0.80
UCI 57 600s ($16$h) 208.5 ± 335.5 200 15.4 ± 31.67 55542.5 ± 114021.12 0.83
MOOC 1200s ($\nicefrac{{1}}{{3}}$h) 199.3 ± 167.2 200 0.3 ± 0.72 1242.8 ± 2597.53 0.88
Wikipedia 3600s ($1$h) 211.7 ± 56.3 200 0.9 ± 0.26 3382.2 ± 921.97 0.89
LastFM 21 600s ($6$h) 204.4 ± 120.2 200 5.9 ± 5.37 21099.5 ± 19331.76 0.91
Myket 5400s ($\nicefrac{{3}}{{2}}$h) 220.1 ± 133.6 200 1.4 ± 1.18 4879.2 ± 4236.54 0.91
Social Evo. 1800s ($\nicefrac{{1}}{{2}}$h) 186.1 ± 165.3 200 0.6 ± 1.26 1984.1 ± 4533.72 0.91
Reddit 900s ($\nicefrac{{1}}{{4}}$h) 226.0 ± 54.5 200 0.2 ± 0.06 792.5 ± 199.35 0.92
The results are presented in Table 3 for continuous-time and Table 4 discrete-time temporal graphs. The tables show AUC-ROC scores for time window-based link forecasting and the relative change compared to the batch-based evaluation commonly used for dynamic link prediction. For continuous-time temporal graphs, the change in performance between our window-based and the batch-based approach largely depends on the dataset: Datasets with a similar window duration for all fixed-sized batches (quantified with high NMI scores in Table 2), such as Wikipedia, Reddit, or Myket, only exhibit small differences between the performances. This is expected since we chose the horizon $h$ to produce batches of the same average size as the fixed-sized batches. Nevertheless, the time windows in datasets with inhomogeneously distributed temporal activity such as Enron or UCI do not fit the fixed-sized batches well (lower NMI in Table 2), showing substantial performance changes across models. This highlights that the performance scores of batch-based evaluation are skewed and may not reflect the models’ performance in a real-world setting on inhomogeneous temporal datasets.
@@
Table 3: Test AUC-ROC scores for link forecasting and the relative change compared to link prediction for continuous-time graphs on the same trained models (standard deviations in Appendix D). We compute the AUC-ROC score per time window and average by weighing each time window equally, regardless of the number of edges (Appendix E discusses additional weighting schemes). The last row/column provides mean $\mu$ and standard deviation $\sigma$ of the absolute relative change per column/row.
@@
Enron 84.0( $\uparrow$8.6%) 80.3( $\uparrow$9.2%) 67.9( $\downarrow$0.2%) 69.0( $\uparrow$17.6%) 75.7( $\uparrow$13.9%) 82.7( $\uparrow$3.6%) 75.1( $\uparrow$11.1%) 88.6( $\uparrow$9.0%) 84.5( $\uparrow$10.6%) 9.3%±5.1%
@@
Myket 64.4( $\uparrow$0.6%) 64.1( $\downarrow$0.1%) 61.2( $\uparrow$0.1%) 57.8( $\uparrow$0.4%) 33.5( $\uparrow$3.1%) 52.6( $\uparrow$1.3%) 58.2( $\downarrow$0.3%) 59.8( $\uparrow$0.5%) 33.8( $\uparrow$3.0%) 1.0%±1.2%
@@
Reddit 80.6( $\downarrow$0.0%) 79.5( $\uparrow$0.0%) 80.4( $\downarrow$0.0%) 78.6( $\downarrow$0.1%) 80.2( $\downarrow$0.0%) 78.6( $\downarrow$0.1%) 76.2( $\downarrow$0.1%) 77.1( $\downarrow$0.1%) 80.2( 0.0%) 0.0%±0.1%
@@
We further observe that, for dynamic link forecasting, the performance of memory-based models on discrete-time temporal graphs tends to decrease more than for other methods. This is expected since these models incorporate information about the present snapshot by updating their memory based on prior batches, which means using part of the snapshot’s edges to predict its remaining edges. Our evaluation method prevents this information leakage, which explains the substantial drop in performance. For Contacts, the discrete-time dataset with the highest NMI score, we see the smallest changes in model performance. This demonstrates that the models’ performance obtained through batch-oriented evaluation reflects the time-window-based performance more closely when a given batch size defines more homogeneous time windows. However, this is often not the case in real-world discrete-time temporal graphs with low granularity and large snapshots.
Table 4: Test AUC-ROC scores as in Table 3 but for discrete-time graphs.
@@
UN Tr. 57.7( $\downarrow$12.8%) 50.3( $\downarrow$20.4%) 54.3( $\downarrow$14.0%) 64.1( $\uparrow$3.9%) 67.6( $\uparrow$4.5%) 85.6( $\downarrow$1.0%) 63.7( $\uparrow$4.5%) 68.6( $\uparrow$3.4%) 70.7( $\uparrow$3.4%) 7.5%±6.6%
@@
Flights 67.4( $\downarrow$3.1%) 66.0( $\downarrow$4.3%) 68.1( $\downarrow$1.0%) 72.6( $\uparrow$0.0%) 65.2( $\uparrow$0.3%) 74.6( $\uparrow$0.0%) 70.6( $\downarrow$0.0%) 70.7( $\downarrow$0.0%) 68.6( $\downarrow$0.5%) 1.0%±1.6%
Cont. 95.6( $\uparrow$0.1%) 94.9( $\downarrow$0.5%) 96.6( $\uparrow$0.5%) 95.9( $\uparrow$0.6%) 86.7( $\uparrow$4.1%) 93.0( $\uparrow$0.9%) 95.7( $\uparrow$1.7%) 95.2( $\uparrow$1.1%) 97.7( $\uparrow$0.6%) 1.1%±1.2%
$\mu\pm\sigma$ 8.3%±10.2% 12.9%±12.2% 12.3%±14.1% 1.5%±1.6% 3.2%±2.7% 0.6%±0.5% 1.5%±1.7% 1.1%±1.2% 3.4%±4.7%

### v2

@@ 3.1 Problems in Batch-based Dynamic Link Prediction
In Figure 2 we use normalised mutual information (NMI) (Vinh et al., 2010) to measure the information loss caused by splitting the temporal edges into batches. NMI quantifies how much information observing one random variable conveys about another random variable (see Appendix B for more information). It takes values between 00, meaning “no information”, and $1$, i.e. “full information”. By treating the index $i$ of each batch $B_{i}$ assigned to each edge $(u,v,t)\in B_{i}$ as one random variable and the associated edge’s timestamp $t$ as the other, we can measure the temporal information that is retained after dividing edges into batches. In this case, an NMI value of $1$ means that we can reconstruct the timestamps of edges correctly from their batch number, and a value of 00 means that batch numbers do not carry any information about timestamps, i.e. total loss of information.
@@
In Figure 2(a), larger batches result in more information loss for continuous-time temporal graphs where timestamps have a high resolution. This is because assigning edges that occur at different times to the same batch discards their temporal ordering; the larger the batch size, the more information is lost. A batch size of $b=1$ preserves most temporal information and maximises NMI because we obtain a bijective mapping between almost all timestamps and batch numbers, except when multiple edges occur simultaneously.
@@
Figure 2(b) shows the batch-size dependent NMI for discrete-time temporal graphs. The “optimal” batch size that retains most temporal information depends on the average number of links per snapshot and, thus, on the characteristics of the data. Too small batch sizes impose an ordering on the edges within the snapshots that is not present in the data while too large batches stretch across snapshots and discard the temporal ordering of edges from different batches.
(a) For continuous-time, the temporal ordering of edges within batches is discarded. With increasing batch size, more edges with different timestamps are assigned to the same time window, resulting in increased information loss.
(b) Small batches for discrete-time temporal graphs impose an ordering amongst the edges within snapshots that is not present in the data, which discards the information that edges of the same snapshot occur simultaneously. The NMI has its maximum near the average snapshot size (cf. Table 1) after which the values decrease again similar to continuous-time datasets.
Figure 2: NMI (y-axis) measures the temporal information loss as a function of different batch sizes (x-axis), where smaller NMI values indicate more information loss.
(2) Information Leakage.
The number of edges inside snapshots of discrete-time graphs is often larger than the batch size (cf. Section C.4). Snapshots are, therefore, divided into multiple batches containing edges with the same timestamp but are evaluated sequentially. The same problem can arise in continuous-time graphs in moments of high activity where the number of edges with a specific timestamp surpasses the batch size. Memory-based models utilise information from previous batches to make predictions for later batches by updating their internal node representations. Due to the issues above, this can include edges from the same snapshot. This leads to an unfair advantage for these models due to the information leakage of edge occurrence patterns in the already processed parts of the snapshot.
(3) Varying Time Window Durations.
In Figure 3, we show the temporal activity for a selection of temporal graphs, i.e. the number of time-stamped edges within a given time interval (see Section C.2 for the remaining datasets). The results show that many real-world temporal graphs have highly inhomogeneous activities. For batch-oriented evaluation, this introduces the issue that each fixed-size batch $B_{i}^{+}$ determines a time window with duration $t_{b\cdot(i+1)-1}-t_{b\cdot i}$, i.e., time windows are shorter during periods with higher activity and longer during periods with lower activity (more details in Section C.3).
Figure 3: Real-world datasets exhibit diverse edge occurrence patterns that are visualised using the edge density across time, i.e., histograms of three continuous- and one discrete-time temporal graphs counting the number of edges per timestamp. Dashed lines divide the datasets into 70% train, 15% validation, and 15% test sets as used in Section 4. The other datasets’ histograms are shown in Section C.2.
(4) Tunable Batch Size.
The varying amount of temporal information (see Figure 2) and the changing window durations (cf. Figure 3, 5, and 6) for different batch sizes highlight that changing the batch size influences the dynamic link prediction task. Effectively, the batch size is a hidden hyperparameter that directly impacts the task’s characteristics (and difficulty). This possibility of manipulating the task by changing the batch size makes the results incomparable between models with different batch sizes.
Consider the Myket dataset (Loghmani & Fazli, 2023) as an example: it contains users $v$ and Android applications $u$, connected at time $t$ when user $v$ installs application $u$. The timestamps are provided in seconds and edges occur roughly every 30s on average (cf. Table 1), making the expected time range for a batch with size $b=2$ approximately 30s. Thus, the task is to predict who installs which applications during this time window. Choosing $b=120$ or $b=2880$ turns the task into a prediction problem for approximately the next hour or day, respectively. For empirical validation, we compare the performance of state-of-the-art TGNNs using two different batch sizes during evaluation in Appendix D and show that the performance can vary substantially between different batch sizes.
3.2 Our Solution: Dynamic Link Forecasting
In real-world applications the prediction time window is inherently connected to the problem at hand, necessitating a task formulation that is chosen carefully for each dataset instead of for each model. To address the aforementioned issues, we propose link forecasting, reformulating link prediction as a task that utilizes a fixed prediction time window with a variable number of edges. Compared to dynamic link prediction with a fixed batch size, this task is both easier and harder: it is easier because it limits the window durations in which the temporal ordering of edges is lost and, additionally, does not introduce artificial temporal orderings that are not present in the data. It is harder because it prevents information leakage and ensures that the task cannot be tuned to fit the model.
Motivation.
@@ 3.2 Our Solution: Dynamic Link Forecasting
The study of temporal information is at the centre of time series forecasting (Benidis et al., 2023) where the goal is to forecast the values of a time series for a given time window in the future. We define our task accordingly and interpret the temporal edges $E$ as $n^{2}$ Boolean time series, each of which takes the value 1 when an edge occurs. Standard multivariate models output a value for each timestamp over a forecasting horizon $h$. In large-scale temporal graphs, it is computationally infeasible to forecast the existence of all $n^{2}$ possible links, thus, only a sample of negative edges is considered instead. In continuous-time dynamic graphs, observations are available at high resolution, e.g. seconds, however, for many practical applications, predicting at lower granularity suffices. For example, it is typically enough to predict whether a customer purchases a certain product within the next day or week. Therefore, we consider forecasting for all timestamps $[t+1,t+h]$ during a time window at once instead of for each of them individually, and define the link forecasting task as follows:
Definition 3.1.
@@
Let $G=(V,E)$ be a temporal graph with node features $H_{V}$ and edge features $H_{E}$. Let $h$ be the time horizon and $W_{i}^{+}:=\{(u,v,t)\mid\exists(u,v,t)\in E:i\cdot h\leq t<(i+1)\cdot h\}$ the set of edges in the $i$-th time window. We use $W_{i}^{-}$ to denote a sample of $|W_{i}^{+}|$ negative edges that are sampled using one of the negative sampling approaches described in Appendix A and do not occur as positive edges in time window $[i\cdot h,(i+1)\cdot h)$. We further use $\hat{E}_{i}=\left\{(u,v,t)\mid\exists(u,v,t)\in E:t<i\cdot h\right\}$ to denote the set of past edges for time window $i$. The goal of dynamic link forecasting is to find a model $f_{\theta}(u,v,t|\hat{E}_{i},H_{V},H_{\hat{E}_{i}})$ with parameters $\theta$ that, for all edges $(u,v,t)\in W_{i}^{+}\cup W_{i}^{-}$ in each time window $i$, forecasts whether $(u,v,t)\in W_{i}^{+}$ or $(u,v,t)\in W_{i}^{-}$.
Crucially, this definition makes the evaluation independent of the batch size $b$ and instead introduces a time horizon $h$ that defines the forecasting time window and coarse-grains the temporal information. This requires deliberately choosing the temporal resolution based on the characteristics of each dataset and application (see Appendix E for examples). Similar to batch-based dynamic link prediction, adopting such a time-window-based perspective enables parallel processing of edges by aggregating temporal information. To summarize, link forecasting is based on a time horizon that affects and controls the amount of temporal information and ensures that the information is consistent for different models, facilitating fair performance comparisons.
Computational Cost
Let $m$ be the total number of links, where typically $m\ll n^{2}$ since real-world graphs are sparse, then we can assign all links to their corresponding time window in a single pass over the links in $O(m)$ by checking their interaction time. After each link is assigned, the time complexity during model evaluation is the same as for the batch-based approach. Since the number of links per time window varies and windows can become large during periods when many temporal edges occur, memory overflows cannot be precluded entirely. To mitigate this issue, time windows can be subdivided into smaller chunks for GPU-based computations. Since information leakage instead of information loss needs to be prevented between chunks of the same time window, steps such as memory updates need to be accumulated and done based on the whole time window. We report the runtime of both approaches for the experiments of Section 4 in Table 11.
@@
We provide implementations for our evaluation procedure in commonly used PyTorch libraries to simplify the adoption of our approach. Specifically, we implement a new DataLoader called SnapshotLoader [^1] https://github.com/M-Lampert/pytorch_geometric/blob/snapshot-loader/torch_geometric/loader/snapshot_loader.py that replaces the widely used TemporalDataLoader in PyTorch Geometric (Fey & Lenssen, 2019). We extend DyGLib (Yu et al., 2023) with a command line argument horizon that can be used in the evaluation pipeline [^2] https://github.com/M-Lampert/DyGLib. The latter was used for the experiments in this work and can be used to reproduce our results.
4 Link Prediction vs. Forecasting
@@
We now experimentally evaluate the performance of nine state-of-the-art models (Kumar et al., 2019; Trivedi et al., 2019; Xu et al., 2020; Rossi et al., 2020; Wang et al., 2021a, b; Poursafaei et al., 2022; Cong et al., 2023; Yu et al., 2023), both for the (conventional) dynamic link prediction task as well as our proposed dynamic link forecasting task. We use implementation and model configurations provided by DyGLib (Yu et al., 2023) (cf. Appendix F) and repeat each experiment five times to obtain averages. We use historical negative sampling (Poursafaei et al., 2022) and train each model using batch-based training and validation with batch size $b=200$ which was found “to be a good trade-off between speed and update granularity” (Rossi et al., 2020) and adopted in later works (Poursafaei et al., 2022; Yu et al., 2023). Afterwards, we evaluate each trained model with weights obtained by batch-based training using both our proposed time-window-based evaluation method and the common batch-based evaluation approach.
@@
Table 2: Average links per window $\overline{|W_{i}^{+}|}$ and standard deviation for horizon $h$ used in the evaluation. We chose $b=200$, i.e. $\overline{|B_{i}^{+}|}=200$, and appropriate horizons $h$ for each dataset to get $\overline{|W_{i}^{+}|}\approx 200$. NMI uses the time window and batch IDs of the test set (cf. Appendix B) to quantify how much the chosen chunks differ between the approaches.
Dataset $h$ $\overline{|W_{i}^{+}|}$ NMI
Enron 172 800s ($48$h) 214.1 ± 274.1 0.80
UCI 57 600s ($16$h) 208.5 ± 335.5 0.83
MOOC 1200s ($\nicefrac{{1}}{{3}}$h) 199.3 ± 167.2 0.88
Wiki. 3600s ($1$h) 211.7 ± 56.3 0.89
LastFM 21 600s ($6$h) 204.4 ± 120.2 0.91
Myket 5400s ($\nicefrac{{3}}{{2}}$h) 220.1 ± 133.6 0.91
Social 1800s ($\nicefrac{{1}}{{2}}$h) 186.1 ± 165.3 0.91
Reddit 900s ($\nicefrac{{1}}{{4}}$h) 226.0 ± 54.5 0.92
Continuous-Time Temporal Graphs.
Following previous work, we set the batch size to $b=200$ and choose the forecasting horizon $h$ such that the resulting time windows contain $200$ edges on average to facilitate comparability with the batch-oriented approach (see Table 2 for the exact values). We quantify the amount of information shared between the two approaches in Table 2 as the NMI score between time window ID and batch ID (cf. Appendix B for a detailed explanation).
@@
Table 3: Test AUC-ROC scores for link forecasting and the relative performance change compared to link prediction for continuous-time graphs using the same trained models (standard deviations in Section G.2). The last row/column provides mean $\mu$ and standard deviation $\sigma$ of the absolute values (without sign) of the relative change per column/row.
@@
Enron 84.0( $\uparrow$8.6%) 80.3( $\uparrow$9.2%) 67.9( $\downarrow$0.2%) 69.0( $\uparrow$17.5%) 75.7( $\uparrow$13.9%) 82.7( $\uparrow$3.6%) 75.1( $\uparrow$11.1%) 88.6( $\uparrow$9.0%) 84.5( $\uparrow$10.6%) 9.3%±5.1%
@@
Myket 64.4( $\uparrow$0.6%) 64.1( $\downarrow$0.1%) 61.2( $\uparrow$0.1%) 57.8( $\uparrow$0.4%) 33.5( $\uparrow$3.1%) 52.6( $\uparrow$1.3%) 58.3( $\downarrow$0.3%) 59.8( $\uparrow$0.5%) 33.8( $\uparrow$3.0%) 1.0%±1.2%
@@
Reddit 80.6( $\downarrow$0.0%) 79.5( $\uparrow$0.0%) 80.4( $\downarrow$0.0%) 78.6( $\downarrow$0.1%) 80.2( $\downarrow$0.0%) 78.6( $\downarrow$0.1%) 76.2( $\downarrow$0.1%) 77.1( $\downarrow$0.1%) 80.2( $\uparrow$0.0%) 0.0%±0.1%
@@
Table 3 shows AUC-ROC scores for time-window-based link forecasting and the relative change compared to the batch-based evaluation of dynamic link prediction (average precision scores are provided in Section G.3). The change in performance between our window-based and the batch-based approach largely depends on the dataset: Datasets with a similar window duration for all fixed-sized batches (quantified by NMI scores close to one in Table 2), such as Wikipedia, Reddit, or Myket, only exhibit small performance differences. This is expected since we chose the horizon $h$ to produce time windows close to the same average size as the fixed-sized batches. Nevertheless, we observe lower NMI values in Table 2 for datasets with inhomogeneously distributed temporal activity such as Enron or UCI – i.e. the time windows do not fit the fixed-sized batches well. These datasets with lower NMIs show substantial performance changes across models. This highlights that the performance scores of batch-based evaluation are skewed and may not reflect the models’ performance in a real-world setting on temporal datasets with inhomogeneous activities.
Figure 4: Model performance of selected continuous-time datasets varies substantially over time which is visualised by the AUC-ROC scores for each batch and time window individually. For both datasets, the best-performing memory-based and non-memory-based model is shown. The time (x-axis) is limited to the test set. The scores for other models and datasets are reported in Section G.4.
To investigate the causes of the performance differences of both evaluation approaches, we analyse the AUC-ROC scores for each batch and time window individually in Figure 4 (and Section G.4 for other datasets and models). We observe that the performance between both approaches diverges most during phases of irregular temporal activity. Specifically, the batch-based evaluation is affected by an anomaly in the Enron dataset at day 1181, where the performance of all models drops substantially. At this time, 1705 temporal edges with identical timestamps are evaluated in 9 consecutive batches. In contrast, our approach uses a single time window for all 1705 edges with identical time stamps. During the second half of the observation period of the Enron dataset, our approach leads to more time windows compared to the number of batches. For our window-based approach, most models perform better during this period compared to their performance in the first half. This increase in performance during the second half of the observation period is masked by the batch-based evaluation, which aggregates this long period of time in a very small number of batches. A similar (daily) pattern can be observed for the MOOC dataset, where each period of high activity is followed by a period of low activity. Our time-window-based evaluation reveals that most models are particularly unreliable during the low activity periods which is again masked by the batch-based approach.
Table 4: Test AUC-ROC scores as in Table 3 for discrete-time graphs.
@@
UN Tr. 57.7( $\downarrow$12.8%) 50.3( $\downarrow$20.4%) 54.3( $\downarrow$14.0%) 64.0( $\uparrow$3.6%) 67.6( $\uparrow$4.5%) 85.6( $\downarrow$1.0%) 63.7( $\uparrow$4.5%) 68.6( $\uparrow$3.4%) 70.7( $\uparrow$3.4%) 7.5%±6.6%
@@
Flights 67.4( $\downarrow$3.1%) 66.0( $\downarrow$4.3%) 68.1( $\downarrow$1.0%) 72.6( $\uparrow$0.0%) 66.0( $\uparrow$0.1%) 74.6( $\uparrow$0.0%) 70.6( $\downarrow$0.0%) 70.7( $\downarrow$0.0%) 68.8( $\uparrow$0.1%) 1.0%±1.6%
Cont. 95.6( $\uparrow$0.1%) 94.9( $\downarrow$0.6%) 96.6( $\uparrow$0.5%) 95.9( $\uparrow$0.6%) 86.7( $\uparrow$4.1%) 93.0( $\uparrow$0.9%) 95.7( $\uparrow$1.3%) 95.2( $\uparrow$1.1%) 97.7( $\uparrow$0.6%) 1.1%±1.2%
$\mu\pm\sigma$ 8.3%±10.2% 12.9%±12.2% 12.3%±14.1% 1.5%±1.5% 3.1%±2.7% 0.6%±0.5% 1.5%±1.7% 1.1%±1.2% 3.4%±4.8%
Discrete-Time Temporal Graphs.
We set $h=1$ to obtain one time window per snapshot, predicting for time intervals ranging from five minutes to a year, depending on the datasets (cf. Table 1). Similar to Table 3, Table 4 shows the AUC-ROC scores and highlights performance differences between both evaluation approaches for discrete-time temporal graphs. For link forecasting, the performance of memory-based models (JODIE, DyRep, TGN) tends to decrease more than for other methods. This is expected since these models incorporate information about the present snapshot by updating their memory based on prior batches, which means using part of the snapshot’s edges to predict its remaining edges — this happens when the batch size is smaller than the snapshot, thus splitting the snapshot into a sequence of batches that are processed sequentially. We designed our evaluation method to prevent this information leakage, which explains the substantial drop in performance. We provide further results showing that our approach prevents information leakage in Section G.5.
Realistic Evaluation.
In the evaluation above, we orient the time horizon $h$ on the batch size $b$ to achieve as much comparability as possible between the link forecasting and prediction tasks. Thereby, we emphasize performance differences between both approaches since we use the same trained models and “only” group the edges differently. In real-world applications, however, the models’ performance should be evaluated using time horizons guided by the specific use case or research question. We provide examples for such an evaluation using the above datasets in Appendix E.